# Large Deviations Principle for Isoperimetry and Its Equivalence to Nonlinear Log-Sobolev Inequalities

### v3

@@ 3.2. Interpretation from Concentration of Measure
The isoperimetric problem with thick boundary admits a natural interpretation from the perspective of concentration of measure. Concentration of measure principle states that for any set of not too small probability, slightly enlarging this set will always have probability close to one. Formally, let $r(a)$ be the minimum number such that the $r$-enlargement of any set of measure $a$ will always have probability close to one. Denote $a=e^{-n\alpha}$ and $r^{2}(a)=n\tau(\alpha)$. By our results, for sufficiently large $n$, $\tau(\alpha)$ is the minimum $\tau$ such that $\psi(\alpha,\tau)=0$, where $\mathsf{C}=\mathsf{W}_{2}^{2}$. Solving it yields $\tau\geq\mathsf{C}(\pi_{X|W},\nu|\pi_{W})$ for all $\pi_{XW}$ such that $D(\pi_{X|W}\|\nu|\pi_{W})\leq\alpha$. Hence, $\tau(\alpha)$ satisfies
$\tau(\alpha)$ $=\sup_{\pi_{XW}:D(\pi_{X|W}\|\nu|\pi_{W})\leq\alpha}\mathsf{C}(\pi_{X|W},\nu|\pi_{W}),$
@@
$\tau_{0}(\alpha):$ $=\sup_{\pi:D(\pi\|\nu)\leq\alpha}\mathsf{C}(\pi,\nu).$
@@
Recall that $K_{C}$ is the optimal transport constant in the following inequality: $\mathsf{W}_{2}(\pi,\nu)\leq\sqrt{\frac{2}{K_{C}}D(\pi\|\nu)},\;\forall\pi.$ Hence, $\tau(\alpha)\leq\frac{2\alpha}{K_{C}},\;\forall\alpha>0,$ and the constant $K_{C}$ cannot be improved. This implies that for sufficiently large $n$, enlarging any set of measure $\exp(-\frac{K_{C}r^{2}}{2})$ by the distance $r$ will always make it have probability close to one, and moreover, the constant $K_{C}$ cannot be improved. In contrast, another (but indeed equivalent) interpretation of $K_{C}$ in the concentration of measure is that for all $A$ such that $\nu_{n}(A)=1/2$, it always holds that $\nu_{n}((A^{r})^{c})\leq\exp(-\frac{K_{C}r^{2}}{2}),\;\forall r>0$ for some constant $C$. Here, the constant $K_{C}$ cannot be improved either.
@@ 4. From Variant Isoperimetry to Standard Isoperimetry
We now turn back to the standard isoperimetric problem in the Riemannian manifold. The “boundary” in the definition of perimeter in this setting is infinitely thin. We now connect the isoperimetric problem with thick “boundary” considered in the last section to this setting.
@@
Bayle [10] and Morgan [49, 51] showed that isoperimetric minimizers always exist and the isoperimetric profile is concave. See an explicit explanation in [45, p. 216].
@@
Under Assumption 2, the following hold.
@@
Theorem 12 (Generalized Heintze-Karcher due to Morgan [49, Theorem 2, Remark 3]).
@@
Note that the sign of the total curvature $H_{S}^{\sigma}(p)$ here is opposite to the one in [49, Theorem 2, Remark 3]. Based on the two theorems above, we build a bridge between the standard isoperimetric problem and the isoperimetric problem with thick boundary.
@@
Assume that Assumption 1 holds. Let $A\subseteq\mathcal{M}^{n}$ denote an isoperimetric minimizer in $(\mathcal{M}^{n},\nu_{n})$ of given measure $a\in(0,1)$ given in Theorem 11. Then, for any $r>0$,
$\frac{I_{n}(a)}{a}=\frac{\nu_{n}^{+}(A)}{\nu_{n}(A)}\geq\frac{1}{r}\ln\frac{\nu_{n}(A^{r})}{\nu_{n}(A)}.$
@@ 5. Putting Everything Together: From Log-Sobolev to Standard Isoperimetry
In this section, we combine all the tools built in Sections 24 to prove the lower bound in our large deviations theorem, i.e., to deduce standard isoperimery from nonlinear log-Sobolev inequalities. To this end, we first deduce variant isoperimery from nonlinear log-Sobolev inequalities by combining Theorem 6 and Theorem 10. Define
$\Upsilon(s,\alpha)=\int_{s}^{\alpha}\frac{1}{\sqrt{\theta(r)}}{\text{\rm d}}r.$
Theorem 14 (From Nonlinear Log-Sobolev to Variant Isoperimetry).
Let $\mathcal{M}$ be a Riemannian manifold equipped with a reference probability measure $\nu=e^{-V}\mathrm{vol}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),V\in C^{2}(\mathcal{M})$. Assume that nonlinear log-Sobolev inequality
$\theta(D(\pi\|\nu))\leq I(\pi\|\nu),\;\forall\pi\in P_{2}^{\mathrm{ac}}(\mathcal{M})$
holds for some convex $\theta$. Assume $\Upsilon(0,\alpha)<\infty$ for some (thus all) $0<\alpha<\infty$. Let $E_{n}$ and $\psi$ be defined in Theorem 10 (for variant isoperimetry) with $\nu=\nu_{X}=\nu_{Y}$ and $p=2$. Then, under Assumption 2, for any $\alpha,\tau\geq 0$,
(5.1) $\lim_{n\to\infty}E_{n}(\alpha,\tau)\leq\psi(\alpha,\tau)$ $\leq\sup_{\begin{subarray}{c}\lambda\in(0,1),\alpha_{0},\alpha_{1}>0:\\ \lambda\alpha_{1}+(1-\lambda)\alpha_{0}\leq\alpha\end{subarray}}\lambda s_{1}^{*}+(1-\lambda)s_{0}^{*},$
where $(s_{0}^{*},s_{1}^{*})=(0,0)$ if
(5.2) $\lambda\Upsilon^{2}(0,\alpha_{1})+(1-\lambda)\Upsilon^{2}(0,\alpha_{0})\leq\tau,$
and otherwise, $(s_{0}^{*},s_{1}^{*})=(s_{0}^{*}(\lambda,\alpha_{0},\alpha_{1}),s_{1}^{*}(\lambda,\alpha_{0},\alpha_{1}))$ is the unique solution to
(5.3) $\Upsilon(s_{0},\alpha_{0})$ $=\sqrt{\frac{\tau\theta(s_{0})}{\lambda\theta(s_{1})+(1-\lambda)\theta(s_{0})}},$
(5.4) $\Upsilon(s_{1},\alpha_{1})$ $=\sqrt{\frac{\tau\theta(s_{1})}{\lambda\theta(s_{1})+(1-\lambda)\theta(s_{0})}}.$
Moreover, the upper bound in (5.1) is strictly increasing in $\alpha$, and hence, the constraint $\lambda\alpha_{1}+(1-\lambda)\alpha_{0}\leq\alpha$ can be replaced by $\lambda\alpha_{1}+(1-\lambda)\alpha_{0}=\alpha$.
@@
Obviously, Assumption 2 implies $\mathbb{E}_{\nu}[d^{q}(X,x_{0})]<\infty$ for any $q\geq 1$. By Theorem 10, the first inequality in (5.1) is obtained. We next prove the second inequality.
Denote $\alpha_{i}=D(\pi_{X|W=i}\|\nu)$, $i\in\{0,1\}$. By Theorem 6, given $\pi_{X|W}$ satisfying $\pi_{X|W=i}\in P_{2}^{\mathrm{ac}}(\mathcal{M}),i\in\{0,1\}$ and given every $s_{i}\leq\alpha_{i}$, there is $\pi_{Y|W}$ such that $\pi_{Y|W=i}\in P_{2}^{\mathrm{ac}}(\mathcal{M})$, $s_{i}=D(\pi_{Y|W=i}\|\nu)$, and
$\mathsf{W}_{2}(\pi_{X|W=i},\pi_{Y|W=i})\leq\Upsilon(s_{i},\alpha_{i}),\;i\in\{0,1\}.$
Hence, it holds that
(5.5) $\psi(\alpha,\tau)$ $\leq\sup_{\begin{subarray}{c}\lambda\in[0,1],\alpha_{0},\alpha_{1}\geq 0:\\ \lambda\alpha_{1}+(1-\lambda)\alpha_{0}\leq\alpha\end{subarray}}\min_{\begin{subarray}{c}s_{0}\in[0,\alpha_{0}],s_{1}\in[0,\alpha_{1}]:\\ \lambda\Upsilon^{2}(s_{1},\alpha_{1})+(1-\lambda)\Upsilon^{2}(s_{0},\alpha_{0})\leq\tau\end{subarray}}\lambda s_{1}+(1-\lambda)s_{0}.$
Observe that when $\alpha$ increases, we can enlarge $s$ to maintain $\Upsilon(s,\alpha)$ unchanged. This implies the RHS above is strictly increasing in $\alpha$.
Observe that $\Upsilon^{2}(s,\alpha)$ is convex in $s$, which implies that given $\lambda\in[0,1],\alpha_{0},\alpha_{1}\geq 0$, the inner minimization in (5.5) is a convex optimization program. We then apply the Lagrangian multiplier method to obtain necessary conditions for optimal solutions. We first assume $\lambda\in(0,1)$. For this case, the optimal solution is $(0,0)$ if (5.2) is satisfied, and otherwise, any optimal solution $(s_{0},s_{1})$ must satisfy
(5.6) $\lambda\Upsilon^{2}(s_{1},\alpha_{1})+(1-\lambda)\Upsilon^{2}(s_{0},\alpha_{0})$ $=\tau,$
(5.7) $\frac{\Upsilon(s_{1},\alpha_{1})}{\sqrt{\theta(s_{1})}}$ $=\frac{\Upsilon(s_{0},\alpha_{0})}{\sqrt{\theta(s_{0})}}.$
These two equations can be rewritten as the ones in (5.3) and (5.4).
We now prove that there is a unique solution to the equations (5.6) and (5.7) (or (5.3) and (5.4)). Observe that given any $\alpha_{i}$, $s\mapsto\frac{\Upsilon(s,\alpha_{i})}{\sqrt{\theta(s)}}$ is decreasing in $s\in(0,\alpha_{i}]$. So, from (5.7), for each $s_{1}$, there is unique $s_{0}=s_{0}^{*}(s_{1})$ satisfying (5.7). Moreover, $s_{0}^{*}(s_{1})$ is increasing in $s_{1}$, which implies that
$s_{1}\mapsto\lambda\Upsilon^{2}(s_{1},\alpha_{1})+(1-\lambda)\Upsilon^{2}(s_{0}^{*}(s_{1}),\alpha_{0})$
is decreasing in $s_{1}\in(0,\alpha_{1}]$. So, there is a unique solution $s_{1}=s_{1}^{*}$ ensuring this function equal to $\tau$, which in turn implies that $(s_{1}^{*},s_{0}^{*}(s_{1}^{*})$) is the unique solution to the equations in (5.6) and (5.7). This yields the second inequality in (5.1) under the assumption $\lambda\in(0,1)$.
If $\lambda=1$, then the optimal $s_{1}^{*}$ for the inner minimization is $s_{1}^{*}=\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha_{1})$, and $s_{0}^{*}$ can be arbitrary, where $\Upsilon_{1}^{-1}$ is the inverse of $\Upsilon$ w.r.t. the first argument when the second is fixed to be $\alpha_{1}$. Similarly, if $\lambda=0$, then the optimal $s_{0}^{*}$ for the inner minimization is $s_{0}^{*}=\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha_{0})$ and $s_{1}^{*}$ can be arbitrary. For both these two cases, the upper bound in (5.5) cannot be greater than $\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha)$. By continuity, one can find that the most right hand side in (5.1) is greater than or equal to $\Upsilon_{1}^{-1}(\sqrt{\tau},\alpha)$. So, the second inequality in (5.1) still holds for $\lambda=0$ or $1$. ∎
We now deduce standard isoperimery from nonlinear log-Sobolev inequalities. Define
(5.8) $\xi(\alpha):=\limsup_{\tau\downarrow 0}\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}},$
where $\psi$ is defined in (3.17) but with $\nu=\nu_{X}=\nu_{Y}$ and $p=2$. As a consequence of Theorem 13 and Theorem 10 (or Theorem 14), we can lower bound the asymptotics of the isoperimetric profile $I_{n}$ in terms of $\xi$ (or $\breve{\Theta}$). Recall that $\breve{\Theta}$ defined in (1.7) is the function characterizing the optimal tradeoff in the nonlinear log-Sobolev inequality. We obtain the lower bound part of Theorem 1. Recall $\underline{\Lambda}$ defined in (1.5).
Theorem 15 (From Nonlinear Log-Sobolev to Standard Isoperimetry).
Under Assumptions 1 and 2, for any $\alpha>0$,
(5.9) $\underline{\Lambda}(\alpha)$ $\geq\xi(\alpha)\geq\sqrt{\breve{\Theta}(\alpha)}.$
Proof.
@@
Let $A_{n}$ be an isoperimetric minimizer of size $e^{-n\alpha}$ in the $n$-dimensional space. From Theorem 13, we see that for any $r>0$,
$\frac{\ln[\nu_{n}(A_{n}^{r})/\nu_{n}(A_{n})]}{r}\leq\frac{\nu_{n}^{+}(A_{n})}{\nu_{n}(A_{n})}=\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}}.$
Setting $r=\sqrt{n\tau}$ and then letting $n\to\infty$, we obtain for any $\tau>0$,
@@
$\underline{\Lambda}(\alpha)$ $\geq\liminf_{n\to\infty}\frac{\ln[\nu_{n}(A_{n}^{\sqrt{n\tau}})/\nu_{n}(A_{n})]}{n\sqrt{\tau}}\geq\frac{\alpha-\psi(\alpha,\tau)}{\sqrt{\tau}},$
where the last inequality follows from Theorem 10. Letting $\tau\to 0$, we obtain the first inequality in (5.9).
We now prove the second inequality in (5.9). If $\underline{\Lambda}(\alpha)=\infty$, then there is nothing to prove. So, we assume $\underline{\Lambda}(\alpha)<\infty$. By the first inequality in (5.9), $\liminf_{\tau\downarrow 0}\psi(\alpha,\tau)=\alpha$. Since $\psi(\alpha,\tau)\leq\alpha$, it holds that $\lim_{\tau\downarrow 0}\psi(\alpha,\tau)=\alpha$. On the other hand, by Theorem 14,
(5.10) $\psi(\alpha,\tau)$ $\leq\sup_{\begin{subarray}{c}\lambda\in(0,1),\alpha_{0},\alpha_{1}>0:\\ \lambda\alpha_{1}+(1-\lambda)\alpha_{0}=\alpha\end{subarray}}\lambda s_{1}^{*}+(1-\lambda)s_{0}^{*},$
where $(s_{0}^{*},s_{1}^{*})$ is $(0,0)$ or the unique solution to (5.3) and (5.4). Therefore, given any $0<\epsilon<\alpha$, for sufficiently small $\tau$, the constraint $\lambda s_{1}^{*}+(1-\lambda)s_{0}^{*}\geq\alpha-\epsilon$ can be posed into the supremization above without changing the optimal value. So, $(s_{0}^{*},s_{1}^{*})=(0,0)$ is excluded. These yield the following bound
$\xi(\alpha)$ $\geq\inf_{\begin{subarray}{c}\lambda\in(0,1),\alpha_{0},\alpha_{1}\geq 0:\\ \lambda\alpha_{1}+(1-\lambda)\alpha_{0}=\alpha,\\ \lambda s_{1}^{*}+(1-\lambda)s_{0}^{*}\geq\alpha-\epsilon\end{subarray}}\frac{\lambda\left(\alpha_{1}-s_{1}^{*}\right)+(1-\lambda)\left(\alpha_{0}-s_{0}^{*}\right)}{\sqrt{\tau}},$
where $(s_{0}^{*},s_{1}^{*})$ is the unique solution to (5.3) and (5.4).
Observe that $\breve{\Theta}$ is increasing. The solution $(s_{0}^{*},s_{1}^{*})$ to (5.3) and (5.4) satisfies
$\frac{\alpha_{1}-s_{1}^{*}}{\sqrt{\breve{\Theta}(s_{1}^{*})}}\geq\int_{s_{1}^{*}}^{\alpha_{1}}\frac{1}{\sqrt{\breve{\Theta}(t)}}{\text{\rm d}}t$ $=\sqrt{\frac{\tau\breve{\Theta}(s_{1}^{*})}{\lambda\breve{\Theta}(s_{1}^{*})+(1-\lambda)\breve{\Theta}(s_{0}^{*})}},$
which implies
$\frac{\alpha_{1}-s_{1}^{*}}{\sqrt{\tau}}$ $\geq\frac{\breve{\Theta}(s_{1}^{*})}{\sqrt{\lambda\breve{\Theta}(s_{1}^{*})+(1-\lambda)\breve{\Theta}(s_{0}^{*})}}.$
Similarly,
$\frac{\alpha_{0}-s_{0}^{*}}{\sqrt{\tau}}\geq\frac{\breve{\Theta}(s_{0}^{*})}{\sqrt{\lambda\breve{\Theta}(s_{1}^{*})+(1-\lambda)\breve{\Theta}(s_{0}^{*})}}.$
Combining these two inequalities and utilizing the convexity of $\breve{\Theta}$ yield that
$\frac{\lambda\left(\alpha_{1}-s_{1}^{*}\right)+(1-\lambda)\left(\alpha_{0}-s_{0}^{*}\right)}{\sqrt{\tau}}$ $\geq\sqrt{\lambda\breve{\Theta}(s_{1}^{*})+(1-\lambda)\breve{\Theta}(s_{0}^{*})}$
$\geq\sqrt{\breve{\Theta}(\lambda s_{1}^{*}+(1-\lambda)s_{0}^{*})}.$
Under the condition $\lambda s_{1}^{*}+(1-\lambda)s_{0}^{*}\geq\alpha-\epsilon$, the last line above is further lower bounded by $\sqrt{\breve{\Theta}(\alpha-\epsilon)}$. Therefore, $\xi(\alpha)\geq\sqrt{\breve{\Theta}(\alpha-\epsilon)}$. Letting $\epsilon\to 0$ yields $\xi(\alpha)\geq\sqrt{\breve{\Theta}(\alpha)}$. ∎

### v4

@@ 3.6. Interpretation from Concentration of Measure
The isoperimetric problem with thick boundary admits a natural interpretation from the perspective of concentration of measure. Concentration of measure principle states that for any set of not too small probability, slightly enlarging this set will always have probability close to one. Formally, let $r(a)$ be the minimum number such that the $r$-enlargement of any set of measure $a$ will always have probability close to one. Denote $a=e^{-n\alpha}$ and $r^{2}(a)=n\tau(\alpha)$. By our results, for sufficiently large $n$, $\tau(\alpha)$ is the minimum $\tau$ such that $\psi(\alpha,\tau)=0$. Solving it yields $\tau\geq\mathsf{W}_{2}(\pi_{X|W},\nu|\pi_{W})$ for all $\pi_{XW}$ such that $D(\pi_{X|W}\|\nu|\pi_{W})\leq\alpha$. Hence, $\tau(\alpha)$ satisfies
$\tau(\alpha)$ $=\sup_{\pi_{XW}:D(\pi_{X|W}\|\nu|\pi_{W})\leq\alpha}\mathsf{W}_{2}(\pi_{X|W},\nu|\pi_{W}),$
@@
$\tau_{0}(\alpha):$ $=\sup_{\pi:D(\pi\|\nu)\leq\alpha}\mathsf{W}_{2}(\pi,\nu).$
@@
Recall that $K_{C}$ is the optimal transport constant in the following inequality: $\mathsf{W}_{2}(\pi,\nu)\leq\sqrt{\frac{2}{K_{C}}D(\pi\|\nu)},\;\forall\pi.$ Hence, $\tau(\alpha)\leq\frac{2\alpha}{K_{C}},\;\forall\alpha>0,$ and the constant $K_{C}$ is optimal. This implies that for sufficiently large $n$, enlarging any set of measure $\exp(-\frac{K_{C}r^{2}}{2})$ by the distance $r$ will always make it have probability close to one, and moreover, the constant $K_{C}$ is the optimal one satisfying this property. In contrast, another (but indeed equivalent) interpretation of $K_{C}$ in the concentration of measure is that for all $A$ such that $\nu_{n}(A)=1/2$, it always holds that $\nu_{n}((A^{r})^{c})\leq\exp(-\frac{K_{C}r^{2}}{2}),\;\forall r>0$. Here, the constant $K_{C}$ cannot be improved either.
@@ 4. From Variant Isoperimetry to Standard Isoperimetry
We now turn back to the standard isoperimetric problem in the Riemannian manifold, in which the boundary is the infinitely thin shell. We now connect the isoperimetric problem with thick shell boundary considered in the last section to this setting.
@@
Bayle [15] and Morgan [55, 56] showed that isoperimetric minimizers always exist and the isoperimetric profile is concave. See an explicit explanation in [52, p. 216].
@@
Under Condition 2, the following hold.
@@
Theorem 11 (Generalized Heintze-Karcher due to Morgan [55, Theorem 2, Remark 3]).
@@
Note that the sign of the total curvature $H_{S}^{\sigma}(p)$ here is opposite to the one in [55, Theorem 2, Remark 3]. Based on the two theorems above, we build a bridge between the standard isoperimetric problem and the isoperimetric problem with thick boundary.
@@
Assume that Condition 1 holds. Let $A\subseteq\mathcal{M}^{n}$ denote an isoperimetric minimizer in $(\mathcal{M}^{n},\nu_{n})$ of given measure $a\in(0,1)$ given in Theorem 10. Then, for any $r>0$,
$\frac{I_{n}(a)}{a}=\frac{\nu_{n}^{+}(A)}{\nu_{n}(A)}\geq\frac{1}{r}\ln\frac{\nu_{n}(A^{r})}{\nu_{n}(A)}\geq\frac{1}{r}\ln\frac{\Gamma_{n}(a,r)}{a}.$
@@ 5. Putting Everything Together: From Log-Sobolev to Standard Isoperimetry
In this section, we combine all the tools built in Sections 24 to prove the lower bound in our large deviations theorem, i.e., to deduce standard isoperimery from nonlinear log-Sobolev inequalities. As a consequence of Theorem 8 and Theorem 12, we can lower bound the asymptotics of the isoperimetric profile $I_{n}$ in terms of $\breve{\Theta}$. Recall that $\breve{\Theta}$ defined in (1.7) is the function characterizing the optimal tradeoff in the nonlinear log-Sobolev inequality. We obtain the lower bound part of Theorem 1. Recall $\underline{\Lambda}$ defined in (1.5).
Theorem 13 (From Nonlinear Log-Sobolev to Standard Isoperimetry).
Under Conditions 1 and 2, for any $\alpha>0$,
(5.1) $\underline{\Lambda}(\alpha)$ $\geq\sqrt{\breve{\Theta}(\alpha)}.$
@@
Let $A_{n}$ be an isoperimetric minimizer of size $e^{-n\alpha}$ in the $n$-dimensional space. From Theorem 12 with $r=\sqrt{n\tau}$, we see that
$\frac{I_{n}(e^{-n\alpha})}{e^{-n\alpha}\sqrt{n}}\geq\frac{\alpha-E_{n}(\alpha,\tau)}{\sqrt{\tau}}.$
@@
Letting $n\to\infty$, from Theorem 8 we obtain for any $\tau>0$, $\underline{\Lambda}(\alpha)\geq\frac{\alpha-\bar{\psi}(\alpha,\tau)}{\sqrt{\tau}}.$ Letting $\tau\to 0$, we obtain $\underline{\Lambda}(\alpha)\geq\limsup_{\tau\downarrow 0}\frac{\alpha-\Upsilon_{\alpha}^{-1}(\sqrt{\tau})}{\sqrt{\tau}}$.
Denote $t=\sqrt{\tau}$ and $s(\alpha,t)=\Upsilon_{\alpha}^{-1}(t)$. Then,
$t=\int_{s(\alpha,t)}^{\alpha}\frac{1}{\sqrt{\breve{\Theta}(r)}}{\text{\rm d}}r.$
Note that $s(\alpha,0)=\alpha$. Taking derivative w.r.t. $t$ at $t=0$ yields $\partial_{t}s(\alpha,t)\big|_{t=0}=\sqrt{\breve{\Theta}(\alpha)}$. Therefore, $\underline{\Lambda}(\alpha)\geq\sqrt{\breve{\Theta}(\alpha)}.$